# Score-based diffusion nowcasting of GOES imagery

### v1

@@ 2.4 Training the neural network
Karras et al. (2022) proved (in their Appendix B.3) that if one has access to an ideal denoiser (i.e., an algorithm that can return an exact copy of a member of the training set given any noised version of that image), then the score function can be computed exactly as
$\nabla_{\bm{x}}\log p\big{(}\bm{x};\sigma(t))=\frac{D(\bm{x};\sigma(t))-\bm{x}}{\sigma(t)^{2}}$ (5)
where the left hand side is the score function and $D(\bm{x};\sigma(t))$ is the ideal denoiser. In practice, we do not generally have access to an ideal denoiser, but denoising is a task that neural networks are already known to be well-suited for, so we can leverage existing techniques to approximate this ideal denoiser. In order to get a good denoiser we train a neural network to minimize the loss from denoising an image, which is typically done with the common $L^{2}$ norm (i.e., Mean Squared Error):
@@
Here $\bm{y}$, are the original non-noised images, $D_{\theta}$ is the neural network denoiser, and $\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ is a multi-dimensional normal distribution with covariance matrix $(\sigma^{2}\mathbf{I})$, where the dimension of the identity matrix, $\mathbf{I}$, is the number of pixels in each image. Thus, $\bm{n}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ yields a matrix (image), $\bm{n}$, with independent Gaussian noise (of mean 00 and variance $\sigma^{2}$) in each pixel. Furthermore, the expectation, $\mathbb{E}$, symbols in Eq. (6) indicate that the neural network is trained to minimize the average taken across all images in the data set and across all noise in the multi-dimensional normal distribution. Know there there is a decent amount of math behind deriving Equation 6, but it is generally derived using the common machine learning process of Evidence Lower BOund (ELBO). Interested readers, please see the appendix of Karras et al. (2022). In practice the loss is also scaled by the current value of $\frac{1}{\sigma}$. The intuition here is that we want to encourage the neural network to perform better at low noise levels (small $\sigma$) because it should be closer to the true solution, so we weight the loss more heavily in that region.
@@
Figure 3 is a visual example of a training step for the denoiser. First, we grab a batch of satellite images ($\bm{y}$). Then, we draw a random noise level for each image in the batch, $\sigma$, from a distribution defined by the user and use that to generate random Gaussian noise images, $\bm{n}$. The noise is then added directly to the images, yielding ($\bm{y}+\bm{n}$), and given to the neural network ($D_{\theta}(\bm{y}+\bm{n};\sigma)$) to be denoised. The choice of neural network architecture is arbitrary so long as it can be good at removing noise from an image. The MSE between the output of the neural network and the original images is then computed and the loss is backwards propagated through the network. This training process is repeated until the loss converges.
Take note of the double expectation in Equation 6 that implies that good sampling of the images, $\bm{y}$, the noise levels, $\sigma$, and of sampled noise images, $\bm{n}$, are required for training. Thus, in order to get a good estimate of the gradient in your learning process, diffusion will likely need a bigger batch than one would expect compared to using a plain U-Net (Ronneberger et al., 2015), because there needs to be a diverse sampling of examples and noise levels.
It is possible to develop a neural network directly to denoise the image, but Karras et al. (2022) developed some clever scalings to make the training task simpler for the neural network. We briefly discuss these scalings here to make it easier to follow the corresponding functions in the code base. Karras et al. (2022) actually chose a neural network, $F_{\theta}(\mbox{input};\mbox{noise level})$, that relates to the denoiser as follows:
@@
where $c_{\text{skip}}(\sigma)$, $c_{\text{out}}(\sigma)$, $c_{\text{in}}(\sigma)$, and $c_{\text{noise}}(\sigma)$ are all scalar functions of $\sigma$ and transform the data such that training and generation are numerically better behaved and thus can be learned more easily by a neural network. The first function, $c_{\text{skip}}(\sigma)$, is not to be confused with traditional skip connections in residual neural networks or U-Nets. Rather, $c_{\text{skip}}(\sigma)$ is a scalar function that provides a scaled version of $\bm{x}$ such that the errors from $F_{\theta}$ are not amplified. The scalar functions $c_{\text{in}}(\sigma)$ and $c_{\text{out}}(\sigma)$ ensure that the input and output of the neural network, $F_{\theta}$, have unit variance. Lastly, $c_{\text{noise}}(\sigma)$ is a scalar function that maps the original noise value so that it can be used better as a condition for the neural network. The derivation of these functions can be found in Appendix B.6 in Karras et al. (2022). The functions are as follows, where $\sigma_{data}$ denotes the scalar variance of the training data, calculated across all pixels:
@@
We found that these functions derived in Karras et al. (2022) worked well for our satellite nowcasting task and suspect that it could be a good starting points also for many Earth and Environmental Science tasks.
@@ 2.5 Generating images with the trained network
The goal of a diffusion model for unconditional image generation is to generate new images that belong to the training dataset from pure Gaussian noise. Thus, to start generating new images we first sample a pure noise image that has a large $\sigma(t)$ value (e.g., 80). In order to translate this pure noise image into an image from the training data we leverage the trained network, $F_{\theta}$, and subsequently, $D_{\theta}$, by substituting the equivalence in Equation 5 into Equation 4
$\mathrm{d}\bm{x}=-\dot{\sigma}(t)~{}\sigma(t)\frac{D_{\theta}(\bm{x};\sigma(t))-\bm{x}}{\sigma(t)^{2}}~{}\mathrm{d}t.$ (12)
Substituting the Karras et al. (2022) choice of $\sigma(t)$ as $t$ yields
$\mathrm{d}\bm{x}=-\frac{D_{\theta}(\bm{x};t)-\bm{x}}{t}~{}\mathrm{d}t.$ (13)
Notice how Equation 13 effectively states that in order to move the pure noisy image from higher noise to lower noise we can use the slope between the denoised image out of the neural network and the original noisy image. Equation 13 is written in its continuous form and like solving the Navier–Stokes equations on a regular grid we can discretize Equation 13
$(\bm{x}_{i}-\bm{x}_{i-1})=-\frac{D_{\theta}(\bm{x}_{i};t_{i})-\bm{x}_{i}}{t_{i}}(t_{i}-t_{i-1})$ (14)
@@
$\bm{x}_{i-1}=\bm{x}_{i}+\big{(}\frac{D_{\theta}(\bm{x}_{i};t_{i})-\bm{x}_{i}}{t_{i}}\big{)}(t_{i}-t_{i-1})$ (15)
The subscript, $i$ denotes the current noisy step, and $(i-1)$ denotes moving backwards towards the original data distribution, i.e., towards the goal.
Figure 4: One step in image generation using score-based diffusion. (a) One step of denoising an image ($\bm{x}_{i}$). The background shading is the same as Figure 2. The circle marker is the location of one pixel within $\bm{x}_{i}$, the star is the denoised value of that same pixel. The transparent image is the result of one step of removing noise. The other white lines are 256 random pixels of $\bm{x}$ and their journey through denoising them. (b) same as in (a) but the trajectories now follow a stochastic trajectory.
Figure 4a shows an example of the image generation starting at $\sigma=3$ (circle marker). The first step is to take the noisy image ($\bm{x}_{i}$) and plug it into $D_{\theta}(\bm{x}_{i};t_{i})$, which gives us the denoised image located at the star. We can now calculate the slope between the circle and star $\left(\frac{D_{\theta}(\bm{x}_{i};t_{i})-\bm{x}_{i}}{t_{i}}\right)$. Plugging into Equation 15 we can take a step of removing noise along this slope and end up at $\bm{x}_{i-1}$ (transparent square and image in Figure 3a). This is repeated until we reach the left hand axis. This is an oversimplified example because the actual markers and lines in Fig. 4 represent changes in individual pixels in the image, $\bm{x}$. Thus, along a single trajectory in an abstract $256\cdot 256$-dimensional space.
Notice that despite updating their slope at each $t_{i}$ value, the lines in Fig. 4a are relatively straight trajectories (faint lines in background). Moreover, because the distribution $p(x;\sigma(t))$ is fixed, these trajectories are completely determined by their origin on the right $y$-axis (i.e., their initial random seed). Recall that we have ignored the second half of Equation 3 in our description so far and that is why these flows are straight and deterministic. In practice, the second part of Equation 3 is not directly implemented in a clean way like the probability flow ODE. Instead, random walks are incorporated by injecting a user defined (i.e., hyperparameter) amount of noise during the denoising procedure. Effectively, think of the same generation procedure as in Fig. 4a, but after we generate $\bm{x}_{i-1}$ we actually add a bit more random noise in (i.e., small step in a random direction towards the right y-axis). This leads the image generation to deviate from the straight ODE paths, and enhances the ensemble image diversity across samples. An example of adding in the stochastic noise during the generation procedure is in Fig. 4b. Trajectories now curve on their journey from the right y-axis to the left y-axis. In the Karras et al. (2022) framework, the scale of the noise added back in is named $S_{churn}$ and the amount of noise is named $S_{noise}$.
2.6 Adding in a condition
@@
The description thus far has described unconditional image generation, that is, generating a random GOES IR image. Many tasks in the Earth and Atmospheric Sciences are much more aligned with conditional image generation. For our paper the task is to generate a specific GOES IR image, the image 10 minutes from now. Most papers on diffusion in the computer science literature tend to build diffusion models for text-to-image tasks (Dall-e , ImageGen; Ramesh et al., 2021; Saharia et al., 2022, respectively). For these text-to-image tasks, incorporating the condition (i.e., the text prompt) is usually done by feeding it into the neural network as a text embedding (i.e., some sort of encoded vector of text). Typically the text embedding relies on known methods to translate the text prompt into tokens and a matrix of data. For our task, image-to-image, it was not initially clear how to include our image as a condition. Through some experimentation we found that including our condition (the past 2 GOES IR images) as additional channels to the input (i.e., concatenate the condition to the noisy dimension) to the denoiser U-Net worked well (Fig. 5 and Fig. 6b). Comparing the results of the unconditional model in Fig. 3 to the results of the conditional model in Fig. 5, we see that the conditional model yields very low errors, while the unconditional model has very large errors, especially for the largest $\sigma$ value, $\sigma=5$. This is not surprising, as the image conditions (bottom right in Fig. 5 provide information about previous time steps, which provide a solid basis for the denoiser to predict the specific weather state we seek to predict, rather than a random weather state.
Furthermore, we found that including the image condition, $\mathbf{c}$, in this manner then the neural network training can easily be expressed in terms of the original Karras framework (Karras et al., 2022) by adding the condition, $\mathbf{c}$, to the score function, which becomes $\nabla_{\bm{x}}\log p\big{(}\bm{x}|\mathbf{c};\sigma(t)\big{)}$, and to the denoiser, which becomes $D_{\theta}(\bm{x}|\mathbf{c};\sigma(t))$. The denoiser is then trained to minimize
@@
This is not the only way to include a condition in diffusion models. Rozet and Louppe (2023), Manshausen et al. (2024), Qu et al. (2024), Martin et al. (2025) and Andry et al. (2025) break the score function into two tasks using Bayes rule:
$\nabla_{\bm{x}}\log p(\bm{x}|\mathrm{\mathbf{c}};\sigma(t))=\underbrace{\nabla_{\bm{x}}\log p(\bm{x};\sigma(t))}_{\text{Unconditional score}}\\ +\underbrace{\nabla_{\bm{x}}\log p(\mathrm{\mathbf{c}}|\bm{x};\sigma(t))}_{\text{Guiding score}}.$ (17)
@@ 3.1 Data
The primary dataset of this paper is Channel 13 (10.3 micron wavelength) from the the Advanced Baseline Imager (ABI; Schmit et al., 2017) on board the GOES-16 satellite, which is part of the Geostationary Operational Environmental Satellite (GOES) satellite series. Channel 13 is a long wave infrared channel and is chosen because of its wide use in general meteorology, particularly for detecting clouds at all times, including both day and night. GOES-16 was launched in 2016 and has been operational since December 2017 located at central longitude of -75 W. Since the motivating goal of this paper is a global cloud forecast, we use full-disk (10 min resolution) data from April 2023 through June 2024.
A full-disk image for Channel 13 has approximately two kilometer resolution near nadir (i.e., over the Amazon) and has a raw image shape of 5,424 by 5,424 pixels. Unfortunately, for the Graphical Processing Units (GPUs) available for this project (a single NVIDIA GH-200 system, hereafter GH-200), the 5,424 by 5,424 image size is prohibitively large and involves gradient computations that would not fit into memory. Thus, we take an approach where we create our datasets by randomly choosing smaller patches of the larger image.
@@ 3.1.4 Preprocessing of images
All images are scaled prior to training the diffusion models such that the mean of the training data is zero and the variance is one. This is critical for diffusion models, because the variance scaling used for the Gaussian noise used during training and inference is highly dependent on the mean and range of the data. For example, if the noise is too small to replace the underlying data distribution, the diffusion models have a challenge generating new data from complete noise. The mean and standard deviation are calculated from the training dataset and applied to all three datasets.

### v2

@@ 2.4 Training the Neural Network
Karras2022 proved (in their Appendix B.3) that if one has access to an ideal denoiser (i.e., an algorithm that can return an exact copy of a member of the training set given any noised version of that image), then the score function can be computed exactly as
$\nabla_{\bm{x}}\log p\big(\bm{x};\sigma)=\frac{D(\bm{x};\sigma)-\bm{x}}{\sigma^{2}}$ (2)
where the left hand side is the score function and $D(\bm{x};\sigma)$ is the ideal denoiser. In practice, we do not generally have access to an ideal denoiser, but denoising is a task that neural networks are already known to be well-suited for, so we can leverage existing techniques to approximate this ideal denoiser. In order to get a good denoiser we train a neural network to minimize the loss from denoising an image, which is typically done with the common $L^{2}$ norm (i.e., Mean Squared Error):
@@
Here $\bm{y}$, are the original non-noised images, $D_{\theta}$ is the neural network denoiser, and $\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ is a multi-dimensional normal distribution with covariance matrix $(\sigma^{2}\mathbf{I})$, where the dimension of the identity matrix, $\mathbf{I}$, is the number of pixels in each image. Thus, $\bm{n}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ yields a matrix (image), $\bm{n}$, with independent Gaussian noise (of mean 0 and variance $\sigma^{2}$) in each pixel. Furthermore, the expectation, $\mathbb{E}$, symbols in Eq. (3) indicate that the neural network is trained to minimize the average taken across all images in the data set and across all noise in the multi-dimensional normal distribution. In practice the loss is also scaled by the current value of $\frac{1}{\sigma}$. The intuition here is that we want to encourage the neural network to perform better at low noise levels (small $\sigma$) because it should be closer to the true solution, so we weight the loss more heavily in that region.
@@
Figure 5 is a visual example of a training step for the denoiser. First, we select a batch of satellite images ($\bm{y}$). Then, we draw a random noise level for each image in the batch, $\sigma$, from a distribution defined by the user and use that to generate random Gaussian noise images, $\bm{n}$. The noise is then added directly to the images, yielding ($\bm{y}+\bm{n}$), and given to the neural network ($D_{\theta}(\bm{y}+\bm{n};\sigma)$) to be denoised. The choice of neural network architecture is arbitrary, as there have been examples of many different architectures (e.g., U-Nets, vision transformers, graph neural networks) that have been used as successful denoisers. The MSE between the output of the neural network and the original images is then computed and the loss is backwards propagated through the network. This training process is repeated until the loss converges.
Take note of the double expectation in Equation 3 that implies that good sampling of the images, $\bm{y}$, the noise levels, $\sigma$, and of sampled noise images, $\bm{n}$, are required for training. Thus, in order to get a good estimate of the gradient in your learning process, diffusion will likely need a bigger batch than one would expect compared to using a plain U-Net (Ronneberger2015), because there needs to be a diverse sampling of examples and noise levels.
The way we have described the training of a denoiser would provide a network sufficient for diffusion, but Karras2022 developed some clever scalings to make the training task simpler for the neural network. We briefly discuss these scalings here to make it easier to follow the corresponding functions in the code base. Karras2022 actually chose a neural network, $F_{\theta}(\mbox{input};\mbox{noise level})$, that relates to the denoiser as follows:
@@
where $c_{\text{skip}}(\sigma)$, $c_{\text{out}}(\sigma)$, $c_{\text{in}}(\sigma)$, and $c_{\text{noise}}(\sigma)$ are all scalar functions of $\sigma$ and transform the data such that training and generation are numerically better behaved and thus can be learned more easily by a neural network. The first function, $c_{\text{skip}}(\sigma)$, is not to be confused with traditional skip connections in residual neural networks or U-Nets. Rather, $c_{\text{skip}}(\sigma)$ is a scalar function that provides a scaled version of $\bm{x}$ such that the errors from $F_{\theta}$ are not amplified. The scalar functions $c_{\text{in}}(\sigma)$ and $c_{\text{out}}(\sigma)$ ensure that the input and output of the neural network, $F_{\theta}$, have unit variance. Lastly, $c_{\text{noise}}(\sigma)$ is a scalar function that maps the original noise value so that it can be used better as a condition for the neural network. The derivation of these functions can be found in Appendix B.6 in Karras2022. The functions are as follows, where $\sigma_{data}$ denotes the scalar variance of the training data, calculated across all pixels:
@@
We found that these functions derived in Karras2022 worked well for our satellite nowcasting task and suspect that it could be a good starting points also for many Earth and Environmental Science tasks.
@@ 2.5 Generating Images with the Trained Network
The goal of a diffusion model for unconditional image generation is to generate new images that belong to the training dataset starting from pure Gaussian noise. Thus, to start generating new images we first sample a pure noise image that has a large $\sigma=\sigma_{\max}$ value (e.g., 80). In order to translate this pure noise image into an image from the training data we leverage the trained network, $F_{\theta}$, and subsequently, $D_{\theta}$, by substituting the equivalence in Equation 2 into Equation 1
$\mathrm{d}\bm{x}=-\sigma\frac{D_{\theta}(\bm{x};\sigma)-\bm{x}}{\sigma^{2}}~\mathrm{d}\sigma.$ (9)
Canceling the $\sigma$ yields
$\mathrm{d}\bm{x}=-\frac{D_{\theta}(\bm{x};\sigma)-\bm{x}}{\sigma}~\mathrm{d}\sigma.$ (10)
Notice how Equation 10 effectively states that in order to move the pure noisy image from higher noise to lower noise we can use the slope between the denoised image out of the neural network and the original noisy image. Equation 10 is written in its continuous form and can be discretized into
$(\bm{x}_{i}-\bm{x}_{i-1})=-\frac{D_{\theta}(\bm{x}_{i};\sigma_{i})-\bm{x}_{i}}{\sigma_{i}}(\sigma_{i}-\sigma_{i-1})$ (11)
@@
$\bm{x}_{i-1}=\bm{x}_{i}+\big(\frac{D_{\theta}(\bm{x}_{i};\sigma_{i})-\bm{x}_{i}}{\sigma_{i}}\big)(\sigma_{i}-\sigma_{i-1})$ (12)
The subscript, $i$ denotes the current noisy step, and (i1)(i-1) denotes moving backwards towards the original data distribution, i.e., towards the goal.
Figure 6: One step in image generation using score-based diffusion. (a) One step of denoising an image ($\bm{x}_{i}$). The background shading is the same as Figure 4 but the image has been rotated to match other literature examples. The circle marker is the location of one pixel within $\bm{x}_{i}$, the star is the denoised value of that same pixel. The transparent image is the result of one step of removing noise. The other white lines are 256 random pixels of $\bm{x}$ and their journey through denoising them. (b) same as in (a) but the trajectories now follow a stochastic trajectory.
Figure 6a shows an example of the image generation starting at $\sigma=3$ (circle marker). The first step is to take the noisy image ($\bm{x}_{i}$) and plug it into $D_{\theta}(\bm{x}_{i};\sigma_{i})$, which gives us the denoised image located at the star. We can now calculate the slope between the circle and star $\left(\frac{D_{\theta}(\bm{x}_{i};\sigma_{i})-\bm{x}_{i}}{\sigma_{i}}\right)$. Plugging into Equation 12 we can take a step of removing noise along this slope and end up at $\bm{x}_{i-1}$ (transparent square and image in Figure 3a). This is repeated until we reach the left hand axis. This is an oversimplified example because the actual markers and lines in Fig. 6 represent changes in individual pixels in the image, $\bm{x}$. Thus, along a single trajectory in an abstract $256\cdot 256$-dimensional space.
Notice that despite updating their slope at each $\sigma_{i}$ value, the lines in Fig. 6a are relatively straight trajectories (faint lines in background). Moreover, because the distribution $p(x;\sigma)$ is fixed, these trajectories are completely determined by their origin on the right $y$-axis (i.e., their initial random seed). To add more randomness and variation to the generation process, random walks are incorporated by injecting a user defined (i.e., hyperparameter) amount of noise during the denoising procedure. Effectively, think of the same generation procedure as in Fig. 6a, but after we generate $\bm{x}_{i-1}$ we actually add a bit more random noise in (i.e., small step in a random direction towards the right y-axis). This leads the image generation to deviate from the straight ODE paths, and enhances the ensemble image diversity across samples. An example of adding in the stochastic noise during the generation procedure is in Fig. 6b. Trajectories now curve on their journey from the right y-axis to the left y-axis. In the Karras2022 framework and the associated code, the scale of the noise added back in is named $S_{churn}$ and the amount of noise is named $S_{noise}$.
2.6 Adding a Condition
@@
The description thus far has described unconditional image generation, that is, generating a random GOES IR image. Many tasks in the Earth and Atmospheric Sciences are much more aligned with conditional image generation. For our paper the task is to generate a specific GOES IR image, the image 10 minutes from now. Most papers on diffusion in the computer science literature tend to build diffusion models for text-to-image tasks (Dall-e , ImageGen; Ramesh2021; Saharia2022, respectively). For these text-to-image tasks, incorporating the condition (i.e., the text prompt) is usually done by feeding it into the neural network as a text embedding (i.e., some sort of encoded vector of text). Typically the text embedding relies on known methods to translate the text prompt into tokens and a matrix of data.
For our task, image-to-image translation, it was not initially clear how to include feature images as a condition. Through some experimentation we found that including our condition (the past 2 GOES IR images) as additional channels to the input (i.e., concatenate the condition to the noisy dimension) to the denoiser U-Net worked well (Fig. 7 and Fig. 8b). Comparing the results of the unconditional model in Fig. 5 to the results of the conditional model in Fig. 7, we see that the conditional model yields very low errors, while the unconditional model has very large errors, especially for the largest $\sigma$ value, $\sigma=5$. This is not surprising, as the image conditions (bottom right in Fig. 7) provide information about previous time steps, which provide a solid basis for the denoiser to predict the specific weather state we seek to predict, rather than a random weather state. Further, this method of adding the conditioning image as an input channel is similar to the conditioning method used in some Conditional Generative Adversarial Networks (CGANs) (e.g., Isola2017).
Including the image condition, $\mathbf{c}$, in this manner can be easily expressed in terms of the original Karras framework (Karras2022) by adding the condition, $\mathbf{c}$, to the score function, which becomes $\nabla_{\bm{x}}\log p\big(\bm{x}|\mathbf{c};\sigma\big)$, and to the denoiser, which becomes $D_{\theta}(\bm{x}|\mathbf{c};\sigma)$. The denoiser is then trained to minimize
@@
This is not the only way to include a condition in diffusion models. Rozet2023, Manshausen2024, Qu_2024, Martin2025 and Andry2025 break the score function into two tasks using Bayes rule:
$\nabla_{\bm{x}}\log p(\bm{x}|\mathrm{\mathbf{c}};\sigma)=\underbrace{\nabla_{\bm{x}}\log p(\bm{x};\sigma)}_{\text{Unconditional score}}\\ +\underbrace{\nabla_{\bm{x}}\log p(\mathrm{\mathbf{c}}|\bm{x};\sigma)}_{\text{Guiding score}}.$ (14)
@@ 3.1 Data
The primary dataset of this paper is Channel 13 (10.3 $\mu m$ wavelength) from the the Advanced Baseline Imager (ABI; schmit2017) on board the GOES-16 satellite, which was part of the Geostationary Operational Environmental Satellite (GOES) satellite series. Channel 13 is a longwave infrared channel and is chosen because of its wide use in general meteorology, particularly for detecting clouds at all times, including both day and night. GOES-16 was operational from December 2017 until April 2025 at a central longitude of -75 W. Since the motivating goal of this paper is a global cloud forecast, we use full-disk (10 min resolution) data from April 2023 through June 2024.
A full-disk image for Channel 13 has approximately two kilometer resolution near nadir (i.e., over the Amazon) and has a raw image shape of 5,424 by 5,424 pixels. Unfortunately, for the Graphical Processing Units (GPUs) available for this project (a single NVIDIA GH-200 system, hereafter GH-200) and the fixed underlying neural network architecture, the 5,424 by 5,424 image size is prohibitively large and involves gradient computations that would not fit into memory. Thus, we take an approach where we create our datasets by randomly choosing smaller patches of the larger image.
@@ 3.1.4 Preprocessing of images
All images are scaled prior to training the diffusion models such that the mean of the training data is zero and the variance is one. While important for all neural networks, it is especially important for diffusion models because the variance scaling used for the Gaussian noise used during training and inference is highly dependent on the mean and range of the data. For example, if the noise is too small to replace the underlying data distribution, the diffusion models have a challenge generating new data from complete noise. The mean and standard deviation are calculated from the training dataset and applied to all three datasets.